The demand equation for a product is

Q= "10*(100-P)^1\/2"

Calculate the consumers’ surplus under equilibrium which occurs at a market price of K84. 


1
Expert's answer
2021-09-13T00:04:38-0400
"Q=10(100-P)^{1\/2}"

"P=100-0.01Q"

"P^*=84=>Q^*=10(100-84)^{1\/2}=40"

The consumer surplus is


"\displaystyle\int_{0}^{Q^*}P(Q)dQ-P^*Q^*"

"=\displaystyle\int_{0}^{40}(100-0.01Q)dQ-84(40)"

"=[100Q-0.005Q^2]\begin{matrix}
 40 \\
 0
\end{matrix}-3360"

"=4000-8-3360=632"



The consumers’ surplus under equilibrium is "632."